# DIAS: A Dataset and Benchmark for Intracranial Artery Segmentation in DSA sequences

### v3

@@ VI-A Impact of Sequences as Inputs
The DSA sequence is a 2D temporal image sequence that represents the contrast agent’s response at different times while flowing through the blood. Theoretically, accurate delineation of vessels requires the aggregation of contrast agent responses from all images in the sequence. In clinical practice, neurosurgeons commonly utilize Minimum Intensity Projection (MIP) to process target sequences in DSA for the visualization of vascular structures [58]. MIP is a straightforward and effective method for aggregating significant vascular information from target sequences. To evaluate the impact of various input modalities on segmentation performance, we conduct comparative experiments that include subsets of the image sequence (First, Last, and First+Last), MIP and average intensity projection (AIP) image as input to the VSS-Net. As shown in Table VI, In ‘First’, where the contrast agent is exclusively displayed in the ICA, the segmentation results are inferior, the DSC is only 62.32%. In comparison, the ‘Last’ as input to the model obtained better results, thanks to the more angiographic information in thin vessel, the DSC is 76.04%. The quantitative results are consistent with those portrayed in Fig. 9 (e) and (f). When using only the first image as the model input, the large vessels are accurately segmented. However, due to the absence of angiographic information for fine vessels in the first image frame, the thin vessel at the arterial termini within the middle cerebral artery (MCA) remain incomplete. Conversely, using only the last image as input enables the accurate segmentation of most small vessels, yet it falls short in precisely delineating the larger vessels in proximity to the ICA.
@@
First image 0.6232 0.9515 0.5388 0.9864 0.4601 0.9049 12.05
Last image 0.7604 0.9647 0.7394 0.9842 0.6146 0.9813 51.89
First + Last image 0.7777 0.9668 0.7589 0.9850 0.6373 0.9846 44.22
Mean 0.7790 0.9663 0.7695 0.9834 0.6398 0.9809 39.25
Min 0.7802 0.9673 0.7631 0.9854 0.6418 0.9842 40.02
Full 0.7822 0.9668 0.7770 0.9838 0.6442 0.9857 43.02
The combination of early and late arterial angiography, known as ‘First+Last’, yielded favorable results, surpassing those of both inputs used independently. As shown in Fig. 9 (a) and (b), although the combination of the ’First’ and ’Last’ images from the arterial phase can display the majority of the vascular structures filled with contrast agent, we can still utilize the intermediate images to supplement details. The experimental findings indicate that employing the complete DSA sequence as input is imperative for achieving optimal performance. The segmentation results of MIP and AIP are quite close to those obtained using the method that utilizes the full DSA sequence as input. Density projection methods such as MIP and AIP convert arterial phase sequences into 2D images along the temporal dimension, which necessitates the pre-selection of target sequences to avoid the influence of capillary and venous phase angiography. Using DSA sequences as the network’s input holds potential in addressing this challenge, as VSS-Net can selectively extract target sequence angiographic information from the entire series, including the capillary and venous phases, for the segmentation of IA.
@@ VII Conclusion
In this paper, we present the DSA-sequence IA segmentation dataset, DIAS, which provides pixel-level annotations along with two forms of weakly supervised annotations. We establish a benchmark for evaluating IA sequence segmentation on DIAS, comprising (1) the Vessel Sequence Segmentation Network, VSS-Net; (2) the Scribble Learning-based Segmentation Framework, SSCR; and (3) the Random Patch-based Self-Training Framework, RPST. Leveraging the DIAS dataset, we have conducted comprehensive experiments to validate the effectiveness of our proposed methodologies, facilitating a deeper understanding of their performance characteristics and offering valuable insights for future research endeavors. Looking ahead, our research will persist, including the expansion of segmentation targets to venous and the segmentation of arteries from complete angiographic sequences. We emphasize that the DIAS dataset and benchmark have been made publicly available, enabling researchers to replicate our work and explore novel algorithms in this domain.

### v4

@@ VI-A Impact of Sequences as Inputs
The DSA sequence is a 2D temporal image sequence that represents the contrast agent’s response at different times while flowing through the blood. Theoretically, accurate delineation of vessels requires the aggregation of contrast agent responses from all images in the sequence. In clinical practice, neurosurgeons commonly utilize Minimum Intensity Projection (MIP) to process target sequences in DSA for the visualization of vascular structures [65]. MIP is a straightforward and effective method for aggregating significant vascular information from target sequences. To evaluate the impact of various input modalities on segmentation performance, we conduct comparative experiments that include subsets of the image sequence (First, Last, and First+Last), MIP and Average Intensity Projection (AIP) image as input to the VSS-Net. As shown in Table VI, In ‘First’, where the contrast agent is exclusively displayed in the ICA, the segmentation results are inferior, the DSC is only 62.32%. In comparison, the ‘Last’ as input to the model obtained better results, thanks to the more angiographic information in thin vessel, the DSC is 76.04%. The quantitative results are consistent with those portrayed in Fig. 10 (e) and (f). When using only the first image as the model input, the large vessels are accurately segmented. However, due to the absence of angiographic information for fine vessels in the first image frame, the thin vessel at the arterial termini within the MCA remain incomplete. Conversely, using only the last image as input enables the accurate segmentation of most small vessels, yet it falls short in precisely delineating the larger vessels in proximity to the ICA.
@@
First 0.6232 0.9515 0.5388 0.9864 0.4601 0.9049 0.6835 12.05
Last 0.7604 0.9647 0.7394 0.9842 0.6146 0.9813 0.6940 51.89
First+Last 0.7777 0.9668 0.7589 0.9850 0.6373 0.9846 0.7039 44.22
Mean 0.7790 0.9663 0.7695 0.9834 0.6398 0.9809 0.6997 39.25
Min 0.7802 0.9673 0.7631 0.9854 0.6418 0.9842 0.7040 40.02
Full 0.7822 0.9668 0.7770 0.9838 0.6442 0.9857 0.7119 43.02
The combination of early and late arterial angiography, known as ‘First+Last’, yielded favorable results, surpassing those of both inputs used independently. As shown in Fig. 10 (a) and (b), although the combination of the ‘First’ and ‘Last’ images from the arterial phase can display the majority of the vascular structures filled with contrast agent, we can still utilize the intermediate images to supplement details. The experimental findings indicate that employing the complete DSA sequence as input is imperative for achieving optimal performance. The segmentation results of MIP and AIP are quite close to those obtained using the method that utilizes the full DSA sequence as input. Density projection methods such as MIP and AIP convert arterial phase sequences into 2D images along the temporal dimension, which necessitates the pre-selection of target sequences to avoid the influence of capillary and venous phase angiography. Using DSA sequences as the network’s input holds potential in addressing this challenge, as VSS-Net can selectively extract target sequence angiographic information from the entire series, including the capillary and venous phases, for the segmentation of IA.
@@ VII Conclusion
In this paper, we present the DSA-sequence IA segmentation dataset, DIAS, which provides pixel-level annotations along with two forms of weakly supervised annotations. We establish a benchmark for evaluating IA sequence segmentation on DIAS dataset, comprising (1) the vessel sequence segmentation network, adept at capturing the spatiotemporal representations in DSA angiography; (2) the weakly-supervised segmentation framework integrating scribble-supervision with consistency regularization; and (3) the random patch-based self-training framework, strategically designed to utilize unannotated DSA data to enhance segmentation performance. Leveraging the DIAS dataset, we have conducted comprehensive experiments to validate the effectiveness of our proposed methodologies, facilitating a deeper understanding of their performance characteristics and offering valuable insights for future research endeavors. Looking ahead, our research will continue, including the expansion of segmentation targets to veins and the segmentation of arteries from complete angiographic sequences. We would like to emphasize that the DIAS dataset and its associated benchmark will be made publicly available, enabling researchers to replicate our work and explore novel algorithms in this domain.